five consecutive odd integers whose sum is 85
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

five consecutive odd integers whose sum is 85
----------------------------------------------------
1st: 2x-3
2nd: 2x-1
3rd: 2x+1
4th: 2x+3
5th: 2x+5
------------------
Equation:
Add to get:
10x +5 = 85
10x = 80
x = 8
----------------
1st: 13
2nd: 15
3rd: 17
4th: 19
5th: 21
====================
Cheers,
Stan H.

Answer by richwmiller(17219)

n= first odd integer
n+2= 2nd
n+4
n+6
n+8
n+n+2+n+4+n+6+n+8=85
5n+20=85
5n=65
n=13
15, 17, 19, 21
No need for all that fancy 2n+1 stuff.
85 is odd so the answer will be odd because we are subtracting 20 from it.